A coffee vending machine is designed to dispense 180 ml of coffee but its owner 

suspects that it is dispensing more than what is designed for. He took a random 

sample of 40 and found out that the mean is 192 ml with a standard deviation of 

4 ml. do you think the owner is right about his suspicion? Test at 0.05 level of 

significance.


1
Expert's answer
2022-05-30T17:44:10-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le180"

"H_1:\mu>180"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=39" and the critical value for a right-tailed test is "t_c =1.684875."

The rejection region for this right-tailed test is "R = \{t:t>1.684875\}."


The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{192-180}{4/\sqrt{40}}=2.1082"


Since it is observed that "t=2.1082>1.684875=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=39" degrees of freedom, "t=2.1082" is "p=0.041491," and since "p= 0.041491<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 180, at the "\alpha = 0.05" significance level.